Ticket Number: HEA-NOC/20020912-01 Ticket Status: OPEN
Ticket Type: Scheduled Ticket Source: HEA-NOC
Ticket Scope: Site Site/Line: HEAnet <==> INEX
Ticket Owner: [log in to unmask] Problem Fixer: HEA-NOC
Ticket Opened: 20020912 08:42:55 UTC Problem Start: 20020912 13:00 UTC+1
Ticket Closed: Problem End:
Change in BGP configuration to accomodate restructuring of peer scheduled for 13:00, 12/9/02.
Traffic to and from HEAnet, particularly for AS5669 and AS9196 may be briefly affected.
Time to Fix:
Estimated 10 minutes, within an hour window.